# Multiscale Hodge Scattering Networks for Data Analysis

### v5

@@ Abstract
We propose new scattering networks for signals measured on simplicial complexes, which we call Multiscale Hodge Scattering Networks (MHSNs). Our construction is based on multiscale basis dictionaries on simplicial complexes, i.e., the $\kappa$-GHWT and $\kappa$-HGLET, which we recently developed for simplices of dimension $\kappa\in\mathbb{N}$ in a given simplicial complex by generalizing the node-based Generalized Haar-Walsh Transform (GHWT) and Hierarchical Graph Laplacian Eigen Transform (HGLET). The $\kappa$-GHWT and the $\kappa$-HGLET both form redundant sets (i.e., dictionaries) of multiscale basis vectors and the corresponding expansion coefficients of a given signal. Our MHSNs use a layered structure analogous to a convolutional neural network (CNN) to cascade the moments of the modulus of the dictionary coefficients. The resulting features are invariant to reordering of the simplices (i.e., node permutation of the underlying graphs). Importantly, the use of multiscale basis dictionaries in our MHSNs admits a natural pooling operation that is akin to local pooling in CNNs, and which may be performed either locally or per-scale. These pooling operations are harder to define in both traditional scattering networks based on Morlet wavelets, and geometric scattering networks based on Diffusion Wavelets. As a result, we are able to extract a rich set of descriptive yet robust features that can be used along with very simple machine learning methods (i.e., logistic regression or support vector machines) to achieve high-accuracy classification systems with far fewer number of parameters to train than most modern graph neural networks. Finally, we demonstrate the usefulness of our MHSNs in three distinct types of problems: signal classification, domain (i.e., graph/simplex) classification, and molecular dynamics prediction.
@@ 1 Introduction
Scattering Transforms were introduced by Mallat in [1] as a method for feature extraction for signals and images. These features are translation quasi-invariant, stable to deformation, and preserve high-frequency information from the input which make them ideal for a wide variety of data classification problems, e.g., texture image classification [2, 3]. In addition, their computational architecture closely resembles a convolutional neural networks (CNNs), which allows for fast, GPU-friendly computation. In fact, these networks are often thought of as a type of CNN, with predetermined wavelet filter banks as their convolution filters and a pointwise modulus operation as their activation function. A key advantage of these networks over traditional CNNs is that since the filter banks do not need to be learned from input data, they are much less data-hungry. Additionally, they are more interpretable since each channel in the hidden representation is a deterministic cascade of wavelet transform convolutions with nonlinear activation and averaging operators.
@@
More recently, Gao et al. introduced an analogous network architecture for node-signals on undirected graphs [4], which they named as “Geometric Scattering (Networks).” In this context, invariance to node-permutation takes the place of translation-in variance. This is achieved in a manner similar to PointNet [5], by aggregating the node features through either a sum or max-values operation into a single value for each channel. This leads to a deformation-stable feature extractor that can be utilized in concert with a very simple learning model (normally logistic regression or support vector machine) to achieve near state-of-the-art classification results on many datasets, with far fewer training examples than CNN-based approaches often require. As a result, these networks generate descriptive yet robust features that can be used along with very simple machine learning methods (i.e., support vector machines or logistic regression) to achieve high-accuracy classification systems with only small amounts of training data and with far fewer number of parameters to train.
@@
In this article, we advance this line of research to apply to signals defined on arbitrarily high-dimensional simplicial structures, i.e., edges, triangles, pyramids, and their $\kappa$-dimensional analogous. Our methods differ from previous work in two key ways. First, previous scattering transform networks have applied only to point-valued signals whereas our construction generalizes to high-dimensional structures. Second, we utilize the $\kappa$-Hierarchical Graph Laplace Transforms ($\kappa$-HGLET) [6, 7] and $\kappa$-Generalized Haar-Walsh Transform ($\kappa$-GHWT)[8, 7] as the wavelet banks in the transforms. Previous work has mostly been based on Morlet wavelets for images and Diffusion Wavelets [9] for graph-based signals. However, we find that the bipartition tree induced by the multiscale transforms we proposed in [7] allows us to form sparser approximations which in turn lead to more efficient feature extraction and therefore more expressive networks. Additionally, the multiscale structure of these bases allows us to easily define local pooling operations which can boost the performance of scattering networks in many applications.
@@ 1.1 Comparison with Related Works
There has been a growth in recent interest in studying signals defined on edges, triangles, and higher-dimensional substructures within graph structured data [10, 11, 12, 13, 14]. Applications in computer vision [15, 16], statistics [17], topological data analysis [14, 18], and network analysis [19] have benefited from the study of high-dimensional simplicial complexes. Convolution-based simplicial neural networks have shown remarkable results in these new domains [20]. We extend this line of research by defining scattering networks on these higher-dimensional domains.
Scattering networks [2] were initially introduced as a tool to explain the success of CNNs on many computer vision problems. These networks had many of the invariant and equivariant properties that make CNNs desirable, but did not contain any learnable ‘filters’, and instead employed a cascade of wavelet convolutions and contractive nonlinearities. Later, Gao et al.  successfully generalized these networks to apply to graph convolutional networks [4]. Our work further generalizes these approaches.
The main ways our MHSNs differ from Geometric Scattering Networks (GSNs) [4] and Deep Haar Scattering Networks (DHSNs) [21] are: 1) MHSNs accept arbitrary simplicial complexes while GSNs/DHSNs were designed for nodes only; and 2) GSNs and DHSNs are based on the Diffusion Wavelets of Coifman and Maggioni [9] and the Haar transform, respectively, and hence they are not based on the hierarchical partitioning of a given graph, while MHSNs are built over the richer HGLET/GHWT dictionaries and more amenable for analysis since they are composed of a collection of orthonormal bases (ONBs). Hodgelets  [22] use a kernel defined in the spectral domain, similar to the spectral graph wavelet transform [23] to define another family of wavelet-like frames for signals on simplicial complexes. Topological Slepians  [24] also form a localized basis dictionary on a given collection of $\kappa$-simplices, but their construction is based on the maximization of primal domain concentration of vectors subject to the dual domain (the frequency domain) support set. However, both Hodgelets and Topological Slepians are difficult to use for scattering transform type representations since they are not hierarchically arranged.
Recently, Chew et al. introduced a method for windowed scattering transforms which achieve local-pooling like operations [25]. However, since the underlying topology of the graph/complex is non-Euclidean, it may be difficult to consistently define the local windows across multiple graphs [26, 27]. It may be possible to use the partitioning scheme proposed in [7] for these windows, but defining appropriate wavelet families for such a hybrid approach requires further study.
@@ 2 Hodge Laplacians and Multiscale Basis Dictionaries
In this section we review some basic Hodge theory to define the Hodge Laplacian on simplicial complexes and then summarize the construction of multiscale basis functions on these spaces. For a more thorough introduction into Hodge Theory see [10, 12, 16] and for a more detailed explanation of multiscale basis dictionaries see [28, 7].
@@ 2.1 Simplicial Complexes and Boundary Operators
[…]
Suppose $\sigma=\{v_{i_{1}},\ldots,v_{i_{\kappa+1}}\}$, $i_{1}<\cdots<i_{\kappa+1}$, and $\alpha\subset\sigma$ is its face. Then, $\sigma\setminus\alpha=\{v_{i_{\ell^{*}}}\}$ for some $\ell^{*}\in\{1,\ldots,\kappa+1\}$. Define the natural parity of $\sigma$ with respect to its face $\alpha$ as $\operatorname{nat}(\sigma,\alpha)\,:=\,(-1)^{\ell^{*}+1}$. An oriented simplex $\sigma$ further has an orientation $p_{\sigma}\in\{\pm 1\}$ which indicates whether its parity with its faces is the same as, or opposite to, its natural parity. When $p_{\sigma}=+1$, we say $\sigma$ is in natural orientation. For example, a directed edge $e=(v_{i},v_{j})$ for $i<j$ is in natural orientation, while if $i>j$, $p_{e}=-1$. An oriented simplicial complex contains at most one orientation for any given simplex.
Given an oriented simplicial complex $C$, for each $\kappa\in\{0,1,\ldots,\kappa_{\mathrm{max}}\}$, the boundary operator is a linear operator ${B_{\kappa}}:X_{\kappa+1}\mapsto X_{\kappa}$, where for $\sigma\in C_{\kappa+1}$, $\alpha\in C_{\kappa}$, the corresponding matrix entries are $[{B_{\kappa}}]_{\alpha\sigma}=p_{\sigma}p_{\alpha}\operatorname{nat}(\sigma,\alpha)$. Likewise, the coboundary operator for each $\kappa\in\{0,1,\ldots,\kappa_{\mathrm{max}}\}$ is just ${B_{\kappa}}^{\scriptscriptstyle{\mathsf{T}}}:X_{\kappa}\rightarrow X_{\kappa+1}$, the adjoint to ${B_{\kappa}}$. The entries of ${B_{\kappa}}$ express relative orientation between simplex and face, and they are a natural way to construct functions taking local signed averages, according to adjacency in the simplicial complex.
@@ 2.2 Hodge Laplacian
The boundary operators just introduced represent discrete differential operators encoding the structure of $\kappa$-regions in a simplicial complex, and so can be building blocks towards a spectral analysis of functions on those regions. For analyzing functions on $\kappa$-simplices with $\kappa>0$, we will construct operators based on the Hodge Laplacian, or $\kappa$-Laplacian. As in [15], the combinatorial $\kappa$-Laplacian is defined for $\kappa$-simplices as
$L_{\kappa}\,:=\,B_{\kappa-1}^{\scriptscriptstyle{\mathsf{T}}}B_{\kappa-1}+B_{\kappa}B_{\kappa}^{\scriptscriptstyle{\mathsf{T}}}\leavevmode\nobreak\ \leavevmode\nobreak\ .$ (1)
Various forms of weighting and normalization are possible, with corresponding advantages and difficulties, and different interpretations of the resulting Hodge Laplacian’s Fiedler vector, as explored in [29, Chap. 4]. In our numerical experiments, we choose the symmetrically normalized, weighted Hodge Laplacian, defined as in [14], as follows. For each $\kappa\in\{0,1,\ldots,\kappa_{\mathrm{max}}\}$, let $D_{\kappa}$ refer to a diagonal matrix, whose diagonal entries contain an assignment of positive weights to each $\kappa$-simplex in $C_{\kappa}$. One such choice, as in [14], is to set a simplex’s weight as its degree, by taking $D_{\kappa+1}=I$, $\operatorname{diag}(D_{\kappa})=|B_{\kappa}|\operatorname{diag}(D_{\kappa+1})$, and $\operatorname{diag}(D_{\kappa-1})=|B_{\kappa-1}|\operatorname{diag}(D_{\kappa})$, where $[|M|]_{\alpha\sigma}=\left|[M]_{\alpha\sigma}\right|$. Define the normalized boundary matrix ${\mathcal{B}}_{\kappa}\,:=\,D_{\kappa}^{-1/2}B_{\kappa}D_{\kappa+1}^{1/2}$. Then the symmetrically normalized, weighted Hodge Laplacian is defined as
$\mathcal{L}_{\kappa}\,:=\,{\mathcal{B}}_{\kappa-1}^{\scriptscriptstyle{\mathsf{T}}}{\mathcal{B}}_{\kappa-1}+{\mathcal{B}}_{\kappa}{\mathcal{B}}_{\kappa}^{\scriptscriptstyle{\mathsf{T}}}\leavevmode\nobreak\ \leavevmode\nobreak\ .$ (2)
Through the rest of this article, when we wish to refer to some variant of the Hodge Laplacian calculated on a $\kappa$-region $R$, without specifying a choice of normalization and/or weighting, we will use $\mathcal{L}(R)$. When $R=C_{\kappa}$, we simplify to $\mathcal{L}\,:=\,\mathcal{L}(C_{\kappa})$.

### v6

@@ Abstract
We propose new scattering networks for signals measured on simplicial complexes, which we call Multiscale Hodge Scattering Networks (MHSNs). Our construction builds on multiscale basis dictionaries on simplicial complexes—namely, the $\kappa$-GHWT and $\kappa$-HGLET—which we recently developed for simplices of dimension $\kappa\in\mathbb{N}$ in a given simplicial complex by generalizing the node-based Generalized Haar–Walsh Transform (GHWT) and Hierarchical Graph Laplacian Eigen Transform (HGLET). Both the $\kappa$-GHWT and the $\kappa$-HGLET form redundant sets (i.e., dictionaries) of multiscale basis vectors and the corresponding expansion coefficients of a given signal. Our MHSNs adopt a layered structure analogous to a convolutional neural network (CNN), cascading the moments of the modulus of the dictionary coefficients. The resulting features are invariant to reordering of the simplices (i.e., node permutation of the underlying graphs). Importantly, the use of multiscale basis dictionaries in our MHSNs admits a natural pooling operation—akin to local pooling in CNNs—that can be performed either locally or per scale. Such pooling operations are more difficult to define in traditional scattering networks based on Morlet wavelets and in geometric scattering networks based on Diffusion Wavelets. As a result, our approach extracts a rich set of descriptive yet robust features that can be combined with simple machine learning models (e.g., logistic regression or support vector machines) to achieve high-accuracy classification with far fewer trainable parameters than most modern graph neural networks require. Finally, we demonstrate the effectiveness of MHSNs on three distinct problem types: signal classification, domain (i.e., graph/simplex) classification, and molecular dynamics prediction.
@@ 1 Introduction
Scattering Transforms were introduced by Mallat in [1] as a method for feature extraction from signals and images. The resulting features are quasi-invariant to translations, stable to certain deformations of the input signals [1, 2], and preserve high-frequency information from the input, making them well-suited for a wide range of data classification tasks, such as texture image classification [3, 4]. In addition, their computational architecture closely resembles a convolutional neural networks (CNNs), which allows for fast, GPU-friendly computation. In fact, these networks are often thought of as a type of CNN, with predetermined wavelet filter banks as their convolution filters and a pointwise modulus operation as their activation function. A key advantage of these networks over traditional CNNs is that since the filter banks do not need to be learned from input data, they are much less data-hungry. Additionally, they are more interpretable since each channel in the hidden representation is a deterministic cascade of wavelet transform convolutions with nonlinear activation and averaging operators.
@@
More recently, Gao et al. introduced an analogous network architecture for node-signals on undirected graphs [5], which they termed “Geometric Scattering (Networks).” In this context, invariance to node permutation takes the place of translation quasi-invariance in the original scattering transform. This is achieved in a manner similar to PointNet [6], by aggregating node features—via either sum pooling or max pooling—into a single value for each channel. The resulting feature extractor is permutation invariant, stable to graph deformations (see also [7, 8]), and can be paired with a simple learning model (typically logistic regression or support vector machine) to achieve near state-of-the-art (SotA) classification results on many datasets, often with far fewer training examples than CNN-based approaches require. As a result, these networks produce descriptive yet robust features that enable high-accuracy classification with small training sets and a minimal number of training parameters.
@@
In this article, we extend this line of research to signals defined on arbitrarily high-dimensional simplicial structures—edges, triangles, pyramids, and their $\kappa$-dimensional analogues. Our methods differ from previous work in two key ways. First, earlier scattering networks have been applied only to point-valued signals, whereas our construction generalizes naturally to higher-dimensional structures. Second, we employ the $\kappa$-Hierarchical Graph Laplace Transform ($\kappa$-HGLET) [9, 10] and the $\kappa$-Generalized Haar–Walsh Transform ($\kappa$-GHWT) [11, 10] as the wavelet filter banks in our transforms. By contrast, most prior work has relied on Morlet wavelets for images and Diffusion Wavelets [12] for graph-based signals. The bipartition tree induced by the multiscale transforms we proposed in [10] enables the construction of sparser approximations, which in turn lead to more efficient feature extraction and, consequently, more expressive networks. Moreover, the multiscale structure of these bases allows us to define local pooling operations in a straightforward manner, which can further enhance the performance of scattering networks in many applications.
@@ 1.1 Comparison with Related Works
There has been growing interest in studying signals defined on edges, triangles, and higher-dimensional substructures within graph-structured data [13, 14, 15, 16, 17]. Applications in computer vision [18, 19], statistics [20], topological data analysis [17, 21], and network analysis [22] have benefited from the study of high-dimensional simplicial complexes. Convolution-based simplicial neural networks have shown remarkable results in these domains [23]. We extend this line of research by defining scattering networks on such higher-dimensional domains.
Scattering networks [3] were initially introduced as a tool to explain the success of CNNs on many computer vision problems. These networks possess many of the invariance and equivariance properties that make CNNs desirable, but they do not contain any learnable filters; instead, they employ a cascade of wavelet convolutions and contractive nonlinearities. Later, Gao et al. successfully generalized these networks to graphs [5]. Our work further generalizes these approaches.
The main ways our MHSNs differ from Geometric Scattering Networks (GSNs) [5] and Deep Haar Scattering Networks (DHSNs) [24] are: a1) MHSNs accept arbitrary simplicial complexes, whereas GSNs and DHSNs were designed for node signals only; and 2) GSNs and DHSNs are based on the Diffusion Wavelets of Coifman and Maggioni [12] and the Haar transform, respectively, and thus are not built on a hierarchical partitioning of the underlying graph. In contrast, MHSNs are constructed over the richer HGLET/GHWT dictionaries, which are more amenable to analysis since they consist of a collection of orthonormal bases (ONBs).
Hodgelets [25] use a kernel defined in the spectral domain—similar to the spectral graph wavelet transform [26]—to define another family of wavelet-like frames for signals on simplicial complexes. Topological Slepians [27] also form a localized basis dictionary on a given collection of $\kappa$-simplices, but their construction is based on maximizing primal-domain concentration of vectors subject to a prescribed dual-domain (frequency-domain) support set. However, both Hodgelets and Topological Slepians are difficult to use for scattering-transform-type representations since they are not hierarchically arranged.
Recently, Chew et al. introduced a method for windowed scattering transforms that achieves local-pooling-like operations [28]. However, because the underlying topology of the graph or complex is non-Euclidean, it may be difficult to consistently define local windows across multiple graphs [29, 30]. It may be possible to use the partitioning scheme proposed in [10] for these windows, but defining appropriate wavelet families for such a hybrid approach requires further study.
@@ 2 Hodge Laplacians and Multiscale Basis Dictionaries
In this section, we review the basic elements of Hodge theory needed to define the Hodge Laplacian on simplicial complexes, and we summarize the construction of multiscale basis functions on these spaces. For a more comprehensive introduction to Hodge theory, see [13, 15, 19]. For a detailed explanation of multiscale basis dictionaries, see [9, 10].
@@ 2.1 Simplicial Complexes and Boundary Operators
In this subsection, we review concepts from algebraic topology to formally define simplicial complexes and introduce several notions of adjacency between simplices. Given a vertex (or node) set $V=\{v_{1},\ldots,v_{n}\}$, a $\kappa$-simplex $\sigma$ is a $(\kappa+1)$-subset of $V$. A face of $\sigma$ is a $\kappa$-subset of $\sigma$, so $\sigma$ has exactly $\kappa+1$ faces. A co-face of $\sigma$ is a $(\kappa+1)$-simplex of which $\sigma$ is a face.
A simplicial complex $C$ is a collection of simplices closed under taking subsets: if $\sigma\in C$ and $\alpha\subset\sigma$, then $\alpha\in C$. In particular, if $\sigma\in C$, all of its faces are also in $C$. Let $\kappa_{\mathrm{max}}(C)\,:=\,\max\left\{\kappa\mid\sigma\in C\ \text{is a $\kappa$-simplex}\right\}$. For each $\kappa\in\{0,1,\ldots,\kappa_{\mathrm{max}}\}$, let $C_{\kappa}$ denote the set of $\kappa$-simplices in $C$, and let $X_{\kappa}$ be the space of real-valued functions on $C_{\kappa}$. When $\kappa>\kappa_{\mathrm{max}}$, we set $C_{\kappa}=\emptyset$. We also refer to $C$ as a $\kappa$-complex when $\kappa_{\mathrm{max}}(C)=\kappa$. A $\kappa$-region of $C$ is any nonempty subset of $C_{\kappa}$.
@@
Figure 1 illustrates these adjacency relations in a toy $2$-complex.
Figure 1: In this small 22-complex $C$, $e_{1}\sim e_{4}$ because they share the face $v_{2}$, and $e_{1}\sim e_{2}$ because they share the face $v_{1}$. Further, $e_{1}\simeq e_{2}$ because their hull $t_{1}\in C$, but $e_{1}\not\simeq e_{4}$, so $e_{1}\>\>\text{\hbox{\set@color\hskip 3.8889pt\hskip-3.8889pt\hbox{\set@color$\sim$}\hskip-3.8889pt\hskip-1.70142pt\raisebox{-3.22221pt}{\hbox{\set@color$\scriptscriptstyle 1$}}\hskip-1.70142pt\hskip 3.8889pt}}\>\>e_{4}$. We have $t_{1}\sim t_{2}$ because they share the face $e_{3}$, and also $t_{1}\>\>\text{\hbox{\set@color\hskip 3.8889pt\hskip-3.8889pt\hbox{\set@color$\sim$}\hskip-3.8889pt\hskip-1.70142pt\raisebox{-3.22221pt}{\hbox{\set@color$\scriptscriptstyle 2$}}\hskip-1.70142pt\hskip 3.8889pt}}\>\>t_{2}$.
Suppose $\sigma=\{v_{i_{1}},\ldots,v_{i_{\kappa+1}}\}$, $i_{1}<\cdots<i_{\kappa+1}$, and let $\alpha\subset\sigma$ be a face. Then $\sigma\setminus\alpha=\{v_{i_{\ell^{*}}}\}$ for some $\ell^{*}\in\{1,\ldots,\kappa+1\}$. The natural parity of $\sigma$ with respect to $\alpha$ is defined as $\operatorname{nat}(\sigma,\alpha)\,:=\,(-1)^{\ell^{*}+1}$. An oriented simplex $\sigma$ has an orientation $p_{\sigma}\in\{\pm 1\}$ indicating whether its parity with its faces agrees with (+1+1) or is opposite to (1-1) its natural parity. When $p_{\sigma}=+1$, we say $\sigma$ is in natural orientation. For example, a directed edge $e=(v_{i},v_{j})$ with $i<j$ is in natural orientation, while if $i>j$, then $p_{e}=-1$. An oriented simplicial complex contains at most one orientation for any given simplex.
Given an oriented simplicial complex $C$, for each $\kappa\in\{0,1,\ldots,\kappa_{\mathrm{max}}\}$, the boundary operator is the linear map ${B_{\kappa}}:X_{\kappa+1}\to X_{\kappa}$, whose matrix entries for $\sigma\in C_{\kappa+1}$ and $\alpha\in C_{\kappa}$ are ${B_{\kappa}}]_{\alpha\sigma}=p_{\sigma}\,p_{\alpha}\,\operatorname{nat}(\sigma,\alpha)$. The coboundary operator for each $\kappa$ is the adjoint ${B_{\kappa}}^{\scriptscriptstyle{\mathsf{T}}}:X_{\kappa}\to X_{\kappa+1}$. The entries of $B_{\kappa}$} encode relative orientation between a simplex and its faces, and naturally define local signed averaging operators according to adjacency in the simplicial complex.
@@ 2.2 Hodge Laplacian
The boundary operators introduced above act as discrete differential operators that encode the structure of $\kappa$-regions in a simplicial complex, and thus serve as building blocks for the spectral analysis of functions on those regions. For analyzing functions on $\kappa$-simplices with $\kappa>0$, we construct operators based on the Hodge Laplacian, or $\kappa$-Laplacian. Following [18], the combinatorial $\kappa$-Laplacian for $\kappa$-simplices is defined as
$L_{\kappa}\,:=\,B_{\kappa-1}^{\scriptscriptstyle{\mathsf{T}}}B_{\kappa-1}+B_{\kappa}B_{\kappa}^{\scriptscriptstyle{\mathsf{T}}}\,.$
Various forms of weighting and normalization are possible, each with its own advantages and challenges, and these choices can also affect how one interprets the Fiedler vector of the resulting Hodge Laplacian; see [31, Chap. 4] for a detailed discussion. In our numerical experiments, we adopt the symmetrically normalized, weighted Hodge Laplacian defined in [17] as follows. For each $\kappa\in\{0,1,\ldots,\kappa_{\mathrm{max}}\}$, let $D_{\kappa}$ be a diagonal matrix whose entries assign positive weights to the $\kappa$-simplices in $C_{\kappa}$. One choice, used in [17], is to set each simplex’s weight to its degree, defined recursively by $D_{\kappa+1}=I$, $\operatorname{diag}(D_{\kappa})=|B_{\kappa}|\operatorname{diag}(D_{\kappa+1})$, and $\operatorname{diag}(D_{\kappa-1})=|B_{\kappa-1}|\operatorname{diag}(D_{\kappa})$, $|M|$ denotes the matrix obtained by taking the absolute value of each entry of $M$, i.e., $|M|\,]_{\alpha\sigma}=\big|[M]_{\alpha\sigma}\big|$. The normalized boundary matrix is then ${\mathcal{B}}_{\kappa}\,:=\,D_{\kappa}^{-1/2}B_{\kappa}D_{\kappa+1}^{1/2}$. The symmetrically normalized, weighted Hodge Laplacian is defined by
$\mathcal{L}_{\kappa}\,:=\,{\mathcal{B}}_{\kappa-1}^{\scriptscriptstyle{\mathsf{T}}}{\mathcal{B}}_{\kappa-1}+{\mathcal{B}}_{\kappa}{\mathcal{B}}_{\kappa}^{\scriptscriptstyle{\mathsf{T}}}\,.$
Throughout the remainder of this article, when referring to a variant of the Hodge Laplacian computed on a $\kappa$-region $R$ without specifying the normalization or weighting, we write $\mathcal{L}(R)$. When $R=C_{\kappa}$, we abbreviate this to $\mathcal{L}\,:=\,\mathcal{L}(C_{\kappa})$.